# Pap2Pat: Towards Automated Paper-to-Patent Drafting using Chunk-based Outline-guided Generation

### v1

@@ 2 Related Work
Table 1: Example for the matching of a PPP. The scores correspond to the respective metrics, the term metrics are shown as min-normalized / max-normalized. Both papers have author lists that contain all the inventors, were published inside the date range, and have titles and abstracts with sufficiently high absolute term similarity to the patent. Since the term similarity scores of paper 1 are higher than those of paper 2 by the specified margin (see Distinctiveness filtering step), paper 1 is correctly matched to the patent.
3 Pap2Pat Dataset
@@
In this section, we present the Pap2Pat dataset containing 1.8k PPPs from a variety of domains, each annotated with multiple outlines. It serves two main purposes: (1) It is an extremely challenging benchmark for LLMs because it contains very long and complex patent documents, and requires deep understanding of the technical domain as well as patent law. (2) It facilitates the development of AI-powered tools for patent drafting, where the patent attorney only performs post-editing rather than writing from scratch, potentially incurring massive cost savings. In the following, we describe the steps taken for constructing the Pap2Pat dataset: scraping and filtering PPPs, parsing the full-text documents and generating the patent outlines.
3.1 Scraping Patent-Paper Pairs
@@
To create Pap2Pat, we start out with the USPTO dataset [^4] https://bulkdata.uspto.gov containing 6.7M patent applications from 2005 to April 2024. For each patent, we query SemOpenAlex (Färber et al. 2023) using SPARQL and retrieve papers with overlapping authors lists and publication dates. Next, we filter the results based on their titles, abstracts, other candidate matches for the same patent, and paper licenses, as elaborated below. Table 2 shows the remaining number of candidates after each filtering step. An example match is shown in Table 1. We perform a systematic manual post-hoc evaluation of the heuristics to validate their precision (see Section 3.2).
Filter Candidates
Authors + Date 930k
+ Term Overlap 100k
+ Distinctiveness 21k
+ Permissive License 1.8k
Table 2: Filter criteria and number of candidates.
Split # pairs # patent tokens # paper tokens # outline bullets
short medium long
train 1000 17808.9 $\pm$ 15132.5 7991.7 $\pm$ 4493.4 36.8 $\pm$ 29.2 73.5 $\pm$ 60.0 149.0 $\pm$ 122.0
val 242 18201.3 $\pm$ 16283.5 8018.5 $\pm$ 4089.2 37.4 $\pm$ 30.8 74.4 $\pm$ 62.9 150.6 $\pm$ 127.5
test 500 18144.9 $\pm$ 13159.5 8081.3 $\pm$ 3919.8 37.5 $\pm$ 24.7 74.9 $\pm$ 50.9 151.6 $\pm$ 103.7
nc-test 71 18415.2 $\pm$ 13995.4 9520.3 $\pm$ 4648.0 37.8 $\pm$ 22.7 76.0 $\pm$ 46.4 154.4 $\pm$ 94.8
all 1813 17977.7 $\pm$ 14724.4 8079.8 $\pm$ 4302.0 37.1 $\pm$ 28.0 74.1 $\pm$ 57.5 150.1 $\pm$ 117.0
Table 3: Dataset Statistics. Values are reported as mean $\pm$ std. Token counts correspond to the Llama-3 tokenizer.
@@
Author Overlap. The patent and the paper of a PPP are by definition authored by overlapping sets of individuals. The overlap of author lists have therefore been identified as an effective (yet not sufficient) criteria for matching PPPs (Magerman, Van Looy, and Song 2010). The requirements for paper authorship are typically much lower than those for patent inventorship (Konski and Wu 2015). In many cases, only the main author(s) and the senior author(s) are listed as inventors. We accordingly employ an asymmetric score that only measures the fraction of inventors $i\in I$ that are also authors $a\in A$, not vice versa:
@@
This score’s effectiveness increases with the number of inventors, so we only consider patents with at least two inventors. Implementating $\text{sim}_{\text{author}}$ requires some form of author name disambiguation to avoid false negatives (different spellings, e.g., with, without, or with abbreviated middle name) and false positives (e.g., very common names like John Smith). There are no author identifiers shared between the patent and paper datasets, so our disambiguation uses the surface names only. To account for false negatives, we use the aliases stored in SemOpenAlex and consider an inventor to be an author if their name matches exactly with one of the aliases. False positives are marginalized by author combinations and subsequent filters: it is highly unlikely that there exist two groups of people with the same set of names working on the same topic at the same time.
@@
Term Overlap. We compare titles and abstracts of patents and papers using term overlap metrics. We first obtain a set of terms $T$ using removal of stopwords and punctuation666based on spaCy’s en_core_web_sm, and stemming777based on NLTK’s PorterStemmer. The score is then computed as the number of shared terms, normalized by the minimum or maximum number of terms, following Magerman, Looy, and Debackere (2015). We additionally weight each term using its IDF value across all candidate titles and abstracts. We compute these IDF weights based on the titles and abstracts of all candidate patents and papers.
$\text{sim}_{\text{term}}(s_{1},s_{2})$ $=\frac{\displaystyle\sum_{t\in T(s_{1})\cap T(s_{2})}\text{idf}(t)}{\displaystyle\text{agg}\Big{(}\sum_{t\in T(s_{1})}\text{idf}(t),\sum_{t\in T(s_{2})}\text{idf}(t)\Big{)}}$
where $s_{1}$ and $s_{2}$ are either the titles or the abstracts of the candidate pair, and $\text{agg}\in\{\min,\max\}$. Thus, we have 4 scores in total, for which we set the thresholds to be 0.15 with $\min$ normalization and 0.1 with $\max$ normalization. We choose the values based on interactive experimentation, but perform a post-hoc validation of the matching precision.
Figure 2: Our approach to chunk-based outline-guided generation. The desired section length information is used to determine chunk sizes for creating outline chunks $o_{1}$, $o_{2}$, …, $o_{n}$.
@@
License. We filter the matches for licenses that allow redistribution and commercial use, i.e., CC-BY, [^8] https://creativecommons.org/licenses/by/4.0/ CC0, [^9] https://creativecommons.org/publicdomain/zero/1.0/ and public domain. We use the license information provided by SemOpenAlex and by the ArXiv API.
3.2 Manual Validation
@@ 3.3 Document Parsing
We parse all patents and papers into a nested JSON schema where each section has a title, paragraphs and subsections field. We make considerable efforts to obtain clean data: we perform LLM-based section hierarchy reconstruction for patents, font-based section hierarchy reconstruction for paper PDFs and formula conversion for patents and paper XMLs from PubMed. We provide more details in Appendix G.
3.4 Outline Generation
One key aspect of our approach are the patent outlines, which simulate the patent attorney’s input to the system in production settings. Examples are shown in Figure 2 and Appendix A. They provide the model with the target headings, as well as bullet points summarizing the document structure and high-level content of every section.
For automatically generating the outlines for Pap2Pat, we leverage the original patents and Llama-3 70B (Dubey et al. 2024). We provide the prompt in Appendix E. To ensure that the model adheres to the output format, we use SGLang (Zheng et al. 2023) for constrained decoding. We enforce a fixed number of bullet points $n_{\text{bullets}}$ per section, where $n_{\text{bullets}}$ is proportional to the length of the text in that section ($n_{\text{chars}}$).
where $l$ is the number of characters that each bullet point summarizes on average. We create outlines in three levels of granularity: long ($l$=500, avg. 150 items), medium ($l$=1000, avg. 74 items) and short ($l$=2000, avg. 37 items), see Table 3. For each section (identified by a heading), we additionally provide the number of characters in the original patent, which can function as the desired content lengths during generation. In a practical application setting, the patent attorney would provide them alongside the outlines.
3.5 Dataset Splits and Statistics
@@
We split our dataset randomly into train (n=1000), test (n=500) and validation (n=242). We additionally create a non-contaminated test set (nc-test) that contains all pairs with a patent published in 2024 (n=71), i.e., after the pretraining cut-off date of all evaluated open-weight LLMs. Thus, we address the concern that LLMs might have seen test data during pretraining (Ravaut et al. 2024). Table 3 shows dataset statistics across the splits, further statistics and plots are presented in Appendix H.
Dataset Split test nc-test
Metric BERTScore ROUGE-L Tokens BERTScore ROUGE-L
Outline short medium long short medium long long long long
Llama-3 8B 67.8 68.3 68.7 23.7 23.0 23.9 51.3% 68.8 22.7
Llama-3 70B 68.6 69.5 70.2 24.7 23.2 24.8 33.6% 70.5 24.1
Mixtral-8x7B 68.1 68.3 69.1 23.6 23.1 25.4 31.7% 68.8 24.3
Qwen2-72B 69.0 69.4 70.2 25.7 24.9 26.1 42.9% 70.4 25.1
Llama-3 8B SFT 69.2 69.9 70.5 26.0 27.2 26.5 162.2% 71.6 26.9
Table 4: Experimental Results. SFT stands for supervised fine-tuned. For the number of tokens and metrics on nc-test, we only show the results regarding the long outline due to space constraints.
4 Chunk-based Outline-guided Generation
In this section, we propose chunk-based outline-guided generation, a novel approach to patent generation illustrated in Figure 2. It serves as a strong baseline for comparison in future work. It is motivated by current LLMs’ limitations regarding sequence length (which needs to cover the instruction, the context, and the output) and therefore generates the patent in chunks, guided by the user-defined outline. We chunk the outline, retrieve paper context based on the outline chunk, and prompt an LLM to generate the patent text for that chunk. Finally, we concatenate the generated outputs and apply only a lightweight post-processing to remove duplicate headings at chunk boundaries.
Chunking and Token Allocation. We assign a fixed number of tokens for the instruction (including system prompt, user prompt and outlines) and equally split the remaining tokens among paper context and output patent. We then chunk the outline using the desired length of the respective patent section. The chunking procedure is designed to pack as much content as possible into each chunk, while preserving the integrity of individual sections. Therefore, we only split a section if it exceeds the token limit on its own.
Retriever. For each chunk $i$, the outline $\text{o}_{i}$ is passed to a retriever, which selects relevant paragraphs from the paper to create the paper context $\text{c}_{i}$. We use BM25 (Robertson and Zaragoza 2009) as the retrieval method and the chunk’s outline as the query. We always include the abstract of the paper because it provides a valuable overview, and successively add paragraphs in the order of their relevance ranking until the token limit is reached. The results are formatted in a Markdown document that comprises all headings and the retrieved paragraphs.
LLMs. The paper context $\text{c}_{i}$, the current outline $\text{o}_{i}$ and prior outlines $\text{o}_{j<i}$ are combined to form the prompt to the LLM (see Appendix F). Based on this prompt, the LLM generates the patent chunk $\text{p}_{i}$, where constrained decoding ensures that the model adheres strictly to the headings in the outline. To make our work reproducible, we only leverage recently published open-weights LLMs that have been reported to achieve state-of-the-art results on other generation tasks. We include a small model (Llama-3 8B (Dubey et al. 2024)) and several larger models (Mixtral-8x7B (Jiang et al. 2024), Llama-3 70B (Dubey et al. 2024) and Qwen2-72B (Yang et al. 2024)). Per default, we use a maximum sequence length of 8k tokens for all LLMs to ensure that the results are comparable and leave the study of the effect of increased maximum sequence length to future work.
Fine-tuning. We fine-tune Llama-3 8B on the training split of Pap2Pat using LoRA (Hu et al. 2022). We use the same prompt and BM25 retrieval as during inference and adopt the hyperparameters proposed by Tribes et al. (2024), see Appendix C.
5 Experiments
We perform a series of experiments using the Pap2Pat dataset to assess the performance of current open-weight LLMs on the proposed task.
5.1 Experimental Setup
To compare the generated and original patents, we use the n-gram-based metric ROUGE-L (F1) (Lin 2004), as well as the neural metric BERTScore (F1) (Zhang et al. 2020) for a balanced evaluation. For BERTScore, we use SciBERT (Beltagy, Lo, and Cohan 2019) as the base model, as it has been shown to be effective in the patent domain (Pujari, Friedrich, and Strötgen 2021). To address the context length limitations of BERT architectures, we process the input in overlapping chunks, where each chunk includes a portion of the preceding chunk to provide contextualization across chunk boundaries. We then compute the dot products between these embeddings across the whole sequence, as in the original variant. The column titled Tokens reports the ratio of the number of generated tokens and number of tokens of the original patent using the Llama-3 tokenizer.
5.2 Results
Table 4 shows the main evaluation results. The three larger models all outperform Llama-3 8B. However, all models generate patents that are much shorter than the original ones, even though the prompt includes the desired length.
BERTScore686868687070707072727272ROUGE-L2323232324242424252525252626262627272727Tokens353535354040404045454545NoPaperAbstractOnlyBM25BM25Oracle
Figure 3: Ablation of retrievers. The plot depicts the results of Qwen2-72B using the long outline.
CS (n=130) Bio (n=152)
BS RL BS RL
Llama-3 8B 67.5 22.1 69.4 25.1
Llama-3 70B 69.0 23.1 70.9 26.2
Mixtral-8x7B 67.5 23.5 70.2 26.7
Qwen2-72B 69.0 24.8 71.0 27.5
Llama-3 8B SFT 70.2 26.2 71.0 26.7
Table 5: Performance comparison between domains (CS=Computer Science, Bio=Biology). The metrics are BERTScore (BS) and ROUGE-L (RL).
Outline Granularity. With all models, the long outline achieves the highest BERTScore and ROUGE-L. We observe a monotonically increasing BERTScore with more detailed outlines. The same is not true for ROUGE-L, where the short outline consistently achieves higher scores than the medium one. Overall, the performance differences between outlines of different granularity are rather small, showcasing the robustness and flexibility of the proposed approach.
Fine-tuning. BERTScore and ROUGE-L improve substantially through fine-tuning, even beyond the scores of the larger models. It is also notable that the generated patents are more than 3 times longer than before fine-tuning. However, the increased length is at least partially due to increased proneness for generating repetitions, which we further discuss in Section 6.
Retrieval. In Figure 3, we show ablation results for the retriever. We include two baselines: NoPaper does not add any context from the paper, and AbstractOnly uses only the abstract. As an upper bound, we use BM25Oracle, where the BM25 query is the original patent text. We find that adding paper context substantially improves performance across metrics, demonstrating that associated papers provide valuable information. We observe a monotonically increasing performance across models and outline granularities. There is a small yet consistent gap between BM25 and BM25Oracle. This suggests that using the outline as a BM25 query is an effective approach, and that more elaborate retrieval methods could close that gap further.
Differences between Domains. We analyze the performance across domains and show the results in Table 5. We include the two most represented domains in the dataset: computer science and biology. All models perform slightly better on patents from biology, with the difference being more pronounced in ROUGE-L than in BERTScore.
Test Data Contamination. We use the non-contaminated test set (nc-test) to study the effect of potential pre-trained memorization on the task performance. If patents are (partially) memorized, one would expect a sudden drop in performance when evaluating on patents published after the pre-training cutoff date. We observe an equal or slightly increased BERTScore and a slightly decreased ROUGE-L, which are more likely attributed to random noise or domain shifts (see Appendix H), and do not indicate systematic issues with LLM memorization affecting our results.
6 Analysis
In this section, we provide deeper insights into the relationship between patents and papers and common error patterns.
@@ 6 Analysis
In general, both patents and papers contain information not present in the other. The paper typically includes more experimental details and insights drawn from the experiments. The patent usually contains more information on the applications and practical benefits of the invention. We analyze the lexical overlaps between the documents and find that only 2.1% of the 4-grams are shared. This underlines the complexity of the task: the two documents describe the same invention from a different perspective using different language. Nevertheless, we find that it is common for attorneys to copy content from the paper to the patent (or vice versa). For instance, many patents and papers share a portion of the figures. But text is also commonly copied between the documents. We find that the longest common substring between paper and patent contains 27 words on average, with a maximum of 495 words. Model-generated patents and corresponding papers, on the other hand, only have a longest common substring of 15 words on average. We furthermore observe that if the paper gives the proposed method or invention a name, the model tends to use that name in the patent as well, whereas the original patents often do not.
The most prevalent error pattern in the generation process is repetition. The models sometimes get stuck in a loop where they repeat the same sentence or paragraph indefinitely. This issue is most pronounced with Llama-3 8B, in particular after fine-tuning, but also occurs with the larger models. We quantify repetitions using RR (Cettolo, Bertoldi, and Federico 2014), a metric that measures the fraction of n-grams that appear just once in a text (lower is better). The patents generated by Llama-3 8B have a much higher RR (53.2 before fine-tuning and 58.9 after fine-tuning), compared to the original patents (42.9) and larger models (Llama-3 70B: 44.1, Qwen2-72B: 40.6, Mixtral-8x7B: 42.3). We hypothesize that this issue is rooted in the inherent repetitiveness of patents. It is common that every paragraph starts with the same words for pages (e.g., ”In some embodiments, component A comprises …”). This is also reflected in the fact that patents have a much higher RR (42.9) than papers (34.5) on our test set. When generating such repetitive text, the model is more likely to get stuck in a loop of repetitions (Holtzman et al. 2020; Fu et al. 2021; Li et al. 2023a). We attempt to alleviate the issue by increasing the temperature and including a repetition penalty. We find that these approaches decrease the number of repetitions but also compromise task performance. Developing specialized methods to avoid exact-repetitions for texts that contain many near-repetitions presents an interesting direction for future research.
In our experiments, we use 8k as chunk size for all models. We also perform preliminary experiments to test whether using a larger chunk size could lead to better and more coherent patents, but find that the repetition problem overshadows these effects. For larger chunk sizes, the repetitions remain uninterrupted for longer, meaning that the content of following outline bullet points does not get addressed. We conclude that the repetition problem must be mitigated before a longer context window can achieve any benefits in automated patent drafting.
7 Conclusion
In this work, we tackle the task of patent generation. Our work is the first to study the generation of patent descriptions and to use PPPs to ground the generation process. We build the Pap2Pat benchmark dataset for training and evaluation, and propose a chunk-based approach to outline-guided paper-to-patent generation and evaluate it in both zero-shot and fine-tuning settings. We find that associated papers provide relevant information, which can be successfully injected into the LLM’s context via a simple retrieval model based on the outline. Furthermore, the approach is robust across different outline granularities and domains. We identify repetitions as the most prevalent error pattern and hypothesize that it is related to the inherent repetitiveness of patents. Promising directions for future research include addressing the repetitions, effectively leveraging models with longer context windows, and utilizing agent frameworks in the generation process.
Limitations
Another limitation of our work is that we have only utilized text overlap metrics, and not utilized more advanced evaluation methods, like with LLMs as judge, or via a user study. While designed to facilitate practical user interaction, the effectiveness of this approach in real-world scenarios remains to be investigated. Future studies will be necessary to fully understand the performance and user experience, and identify areas for improvement.
Ethics Statement
The development of AI-powered patent generation tools may raise ethical concerns. We emphasize that our research is intended to support and augment the work of patent professionals, rather than to automate the patenting process or facilitate malicious activities such as patent trolling.
Appendix A Patent Outline Example

### v3

@@ Appendix A Pap2Pat Dataset
Table 4: Example for the matching of a PPP. The scores correspond to the respective metrics, the term metrics are shown as $\min$-normalized / $\max$-normalized. Both papers have author lists that contain all the inventors, were published inside the date range, and have titles and abstracts with sufficiently high absolute term similarity to the patent. Since the term similarity scores of paper 1 are higher than those of paper 2 by the specified margin (see Distinctiveness filtering step), paper 1 is correctly matched to the patent.
@@
We present the Pap2Pat dataset containing 1.8k PPPs from a variety of domains, each annotated with multiple outlines. It serves two main purposes. First, it is a challenging benchmark for LLMs that requires long-text generation capabilities and deep understanding of the technical domain as well as patent law. Second, it facilitates the development of AI-powered tools for patent drafting, where the patent attorney only performs post-editing rather than writing from scratch, potentially incurring massive cost savings. In the following, we describe the steps taken for constructing the Pap2Pat dataset: scraping and filtering PPPs, parsing the full-text documents and generating the patent outlines.
A.1 Scraping Patent-Paper Pairs
@@
The patent and paper in a PPP typically do not cite each other, so we cannot rely on front-page or in-text citations Marx and Fuegi (2020, 2022) to find PPPs. The matching must rely on document metadata and content alone. We start out with the USPTO dataset [^7] https://bulkdata.uspto.gov containing 6.7M patent applications from 2005 to April 2024. For each patent, we query SemOpenAlex Färber et al. (2023) using SPARQL and retrieve papers with overlapping authors lists and publication dates. Next, we filter the results based on their titles, abstracts, other candidate matches for the same patent, and paper licenses, as elaborated below. Table 2 shows the remaining number of candidates after each filtering step. As can be seen, the major limiting factor for our benchmark size is the issue of scientific articles being published under restrictive licenses. An example match is shown in Table 4. We perform a systematic manual post-hoc evaluation of the heuristics to validate their precision (see Section A.2).
@@
Author Overlap. The patent and the paper of a PPP are by definition authored by overlapping sets of individuals. The overlap of author lists have therefore been identified as an effective (yet not sufficient) criteria for matching PPPs Magerman et al. (2010). The requirements for paper authorship are typically much lower than those for patent inventorship Konski and Wu (2015). In many cases, only the main author(s) and the senior author(s) are listed as inventors. We accordingly employ an asymmetric score that only measures the fraction of inventors $i\in I$ that are also authors $a\in A$, not vice versa:
@@
This score’s effectiveness increases with the number of inventors, so we only consider patents with at least two inventors. Implementing $\text{sim}_{\text{author}}$ requires some form of author name disambiguation to avoid false negatives (different spellings, e.g., with, without, or with abbreviated middle name) and false positives (e.g., very common names like John Smith). There are no author identifiers shared between the patent and paper datasets, so our disambiguation uses the surface names only. To account for false negatives, we use the aliases stored in SemOpenAlex and consider an inventor to be an author if their name matches exactly with one of the aliases. False positives are marginalized by author combinations and subsequent filters: it is highly unlikely that there exist two groups of people with the same set of names working on the same topic at the same time.
@@
Term Overlap. We compare titles and abstracts of patents and papers using term overlap metrics. We first obtain a set of terms $T$ using removal of stopwords and punctuation999based on spaCy’s en_core_web_sm and stemming.101010based on NLTK’s PorterStemmer The score is then computed as the number of shared terms, normalized by the minimum or maximum number of terms, following Magerman et al. (2015). We additionally weight each term using its IDF value across all titles and abstracts of patents and papers.
$\text{sim}_{\text{term}}(pat,pap)$ $=\frac{\displaystyle\sum_{t\in T(pat)\cap T(pap)}\text{idf}(t)}{\displaystyle\text{agg}\Big{(}\sum_{t\in T(pat)}\text{idf}(t),\sum_{t\in T(pap)}\text{idf}(t)\Big{)}}$
where $pat$ and $pap$ are either the titles or the abstracts of the documents, and $\text{agg}\in\{\min,\max\}$. Thus, we have 4 scores in total, for which we set the thresholds to be 0.15 with $\min$ normalization and 0.1 with $\max$ normalization. We choose the values based on interactive experimentation, but perform a post-hoc validation of the matching precision.
@@
License. We filter the matches for licenses that allow redistribution and commercial use, i.e., CC-BY,CC0,and public domain. We use the license information provided by SemOpenAlex and by the ArXiv API.
A.2 Manual Validation
@@ A.3 Document Parsing
We parse all patents and papers into a nested JSON schema where each section has a title, paragraphs and subsections field. We make considerable efforts to obtain clean data: we perform LLM-based section hierarchy reconstruction for patents, font-based section hierarchy reconstruction for paper PDFs and formula conversion for patents and papers. We provide more details in Appendix B.
A.4 Dataset Characteristics
@@
We split our dataset randomly into train (n=1000), test (n=500) and validation (n=242). We additionally create a non-contaminated test set (nc-test) that contains all pairs with a patent published in 2024 (n=71), i.e., after the pretraining cut-off date of all evaluated open-weight LLMs. Thus, we address the concern that LLMs might have seen test data during pretraining Ravaut et al. (2024). Table 1 shows dataset statistics across the splits. Appendix K shows further statistics and plots, including the distribution over domains and the number of pairs over time.
@@
In general, both patents and papers contain information not present in the other. The paper typically includes more experimental details and insights drawn from the experiments. The patent usually contains more information on the applications and practical benefits of the invention. We analyze the lexical overlaps between the documents and find that only 2.1% of the 4-grams are shared. This underlines the complexity of the task: the two documents describe the same invention from a different perspective using different language. Nevertheless, we find that it is common for attorneys to copy content from the paper to the patent (or vice versa). For instance, many patents and papers share a portion of the figures, as well as verbatim copied text.
Appendix B Document Parsing Pipeline
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Appendix C Repetition Removal
Algorithm 1 Repetition Detection
1:function matches(l1, l2)
2:     $n\_match\leftarrow 0$
3:     for $i\leftarrow 0$ to $\text{len}(l1)-1$ do
4:         if $l1[i]=l2[i$then
5:              $n\_match\leftarrow n\_match+1$
6:         end if
7:     end for
8:     return n_match / len(l1) $>$ 0.9
9:end function
10:
11:function detect_repetitions(words, min_length, max_cycle_length)
12:     $n\leftarrow len(words)$
13:     for $k=1$ to max_cycle_length do
14:         if matches(words[-k:], words[-2 * k : -k]) then $\triangleright$ Check for repetition of length k
15:              $i\leftarrow 2$triangleright Number of times the pattern is repeated
16:              while matches(words[-k:], words[-(i + 1) * k : -i * k]) do $\triangleright$ Search for additional occurrences
17:                  $i\leftarrow i+1$
18:              end while
19:              remove_indices $\leftarrow$ [n - (i - 1) * k, …, n]
20:              total_length $\leftarrow i*k$
21:              if total_length $\geq$ min_length then
22:                  words $\leftarrow$ words[: n - (i - 1) * k] $\triangleright$ Keep only first occurrence of pattern
23:                  remove_rest $\leftarrow$ detect_repetitions(words, min_length, max_cycle_length) $\triangleright$ Recursive call for remaining text
24:                  return remove_indices + remove_rest $\triangleright$ Return indices to remove
25:              end if
26:         end if
27:     end for
28:     return []
29:end function
We design a procedure to remove infinite repetitions from generated outputs to study their effect on evaluation metrics. In that context, we characterize an infinite repetition as a sequence of tokens that appears multiple times until the end of the generation. To find such repetitions, we apply the following recursive algorithm in Algorithm 1 to each chunk and remove the returned indices. To account for reptitions where the model alters the patterns slightly (e.g., incrementing a number) in each iteration, we consider two word sequences equal if 90% of their positions are equal. By default, we use min_length = 50 and max_cycle_length = 300.
Appendix D Contamination Results
Tokens Text Sim $\uparrow$ Content-Level Metrics (SCALE) $\uparrow$ Language $\uparrow$ Repetitions
Coverage Factuality
BS R-L GenRef RefGen Ref + PapGen Style DiscoScore RR RR>80
Heuristic Baselines / Skylines
Reference Patent 18.4k (100.0%) 100.0 100.0 88.6 $\pm$ .57 88.5 $\pm$ .23 88.7 $\pm$ .21 99.8 100.0 15.2 0.1
Similar Patent 26.1k (141.8%) 66.0 35.3 31.4 $\pm$ .93 28.0 $\pm$ .64 28.3 $\pm$ .68 53.9 98.1 12.5 0.1
Outline 1.4k (7.6%) 56.7 10.2 42.1 $\pm$ .94 63.4 $\pm$ .88 63.8 $\pm$ .86 25.9 85.1 19.4 0.0
Paper 9.5k (51.7%) 69.6 42.2 46.7 $\pm$ 1.01 46.5 $\pm$ .87 88.8 $\pm$ .41 37.2 98.0 8.2 0.0
[…]
Mixtral-8x7B 3.0k (16.3%) 66.5 21.7 40.7 $\pm$ 1.16 67.2 $\pm$ 1.19 73.9 $\pm$ 1.16 36.5 96.8 15.9 0.1
Qwen2-72B 2.9k (15.5%) 66.4 21.2 42.0 $\pm$ 1.21 65.0 $\pm$ .76 71.9 $\pm$ .23 34.5 96.9 8.9 0.0
   w/o Paper 3.4k (18.2%) 65.9 21.8 41.4 $\pm$ 1.25 66.3 $\pm$ 1.04 67.0 $\pm$ .99 34.1 96.1 9.0 0.2
   w/o Outline 2.0k (11.0%) 63.7 15.8 36.7 $\pm$ 1.27 56.5 $\pm$ .67 75.7 $\pm$ 1.13 30.3 96.8 8.2 0.0
[…]
Llama-3 8B 9.6k (52.1%) 68.8 41.4 42.3 $\pm$ 1.11 60.3 $\pm$ 1.60 65.6 $\pm$ 1.23 36.8 96.6 25.4 2.7
Llama-3 8B SFT 27.0k (146.8%) 71.2 44.0 44.0 $\pm$ 1.47 49.4 $\pm$ 1.42 52.1 $\pm$ 1.36 45.0 97.8 53.7 27.6
   w/ rep. removal 18.2k (99.1%) 72.1 51.1 44.7 $\pm$ 1.03 52.0 $\pm$ 1.13 55.4 $\pm$ 1.15 50.7 98.2 39.6 8.3
Mixtral-8x7B 6.3k (34.4%) 68.8 34.4 45.1 $\pm$ 2.00 64.4 $\pm$ 1.40 70.6 $\pm$ 1.24 39.5 96.9 14.4 0.1
Llama-3 70B 6.1k (33.4%) 70.5 39.5 45.9 $\pm$ 1.15 64.7 $\pm$ 1.06 68.7 $\pm$ .87 44.6 97.0 17.2 0.1
Qwen2-72B 8.2k (44.7%) 70.4 40.1 46.9 $\pm$ 1.26 65.3 $\pm$ .83 69.8 $\pm$ .42 44.3 97.0 10.9 0.0
[…]
Qwen2-72B 17.2k (93.2%) 71.5 50.5 49.9 $\pm$ 1.70 59.2 $\pm$ 1.00 64.9 $\pm$ .43 41.9 96.6 8.8 0.2
Table 5: Experimental Results on the non-contaminated test set. Italic values represent upper bounds that used test data in the prediction. The best value per column is bold, the best per section bold italics. Tokens are reported as absolute and relative to the reference. BS = BERTScore. R-L = ROUGE-L. Text Sim = Standard Text Similarity Metrics.
Appendix E Patent Outline Example